OpenStudy (anonymous):

Solve the following differential equation : sin x dy/dx + cos x.y = cos x.sin2 x

OpenStudy (zzr0ck3r):

what are you doing in class? Are you solving with eigen values? laplace transforms?undetermined coefficients?,,,

OpenStudy (anonymous):

what u wana say @zzr0ck3r

OpenStudy (anonymous):

divide the equation throughout by sinx

OpenStudy (anonymous):

\[\sin(x)*dy/dx + \cos(xy) = \cos(x)*\sin^2(x)\]

OpenStudy (anonymous):

correct?

OpenStudy (calculusfunctions):

No, I think it's double angle, not sin squared. That's what I think. The problem is that people don't always type accurately.

OpenStudy (anonymous):

well then no point in me trying on this one.... Gotta be explicit people!

OpenStudy (calculusfunctions):

I agree. I think it's best to leave it alone until the person comes back on line. Til then time to move on to new adventures.

OpenStudy (anonymous):

there is another way to do it let me show you

OpenStudy (calculusfunctions):

@shruti if you mean\[\sin x \frac{ dy }{ dx }+y \cos x =\cos x \sin 2x\]then this is a first-order linear equation and can now easily be solved. If and when we're both on line, thae I can teach you if you need it. I'll be here for about thirty more minutes right now.

OpenStudy (unklerhaukus):

\[\sin (x) \frac{\text dy}{\text dx} + y\cos (x) = \cos (x)\sin^2 (x)\] \[ \frac{\text dy}{\text dx} + y\cot (x) = \cos (x)\sin (x)\]

OpenStudy (unklerhaukus):

@shruti, please clarify the question, if by sin2 x you mean sin^2 x, then i have a ful solution we can work through